How to do work 3/5 divided by 2/5
Answer by richwmiller(17219)

multiplying by the inverse is the same as division
for example
10 divided by 5 is the same as multiplying by 1/5
so
3/5 divided by 2/5 is the same as multiplying 3/5 *5/2
3/5 *5/2=(3*5)/(2*5)=3/2